OpenStudy (anonymous):

In a city school, 60% of students have blue eyes, 55% have dark hair, and 35% have blue eyes and dark hair. What is the probability (rounded to the nearest whole percent) that a randomly selected student has dark hair, given that the student has blue eyes? 58% 20% 80% 64%

OpenStudy (anonymous):

Oh hello there

OpenStudy (anonymous):

Hi lol

OpenStudy (anonymous):

Can anyone help?

OpenStudy (anonymous):

I dont know how. Give me a few minutes

OpenStudy (anonymous):

OK

OpenStudy (er.mohd.amir):

given that P(blue eye)=P(A)=60%=0.6 P(dark hair)=P(B)=55%=0.55 P(A intersection B)=35%=0.35

OpenStudy (er.mohd.amir):

this problem of conditional probabilty

OpenStudy (anonymous):

Ok one sec

OpenStudy (anonymous):

How do I go about solving this though? What goes where? How do I set it up?

OpenStudy (er.mohd.amir):

P(B/A)=P(A interscetion B)/P(A)

OpenStudy (er.mohd.amir):

simply put values

OpenStudy (anonymous):

Snufflebug02 do you still need explanation

OpenStudy (anonymous):

Kind of I don't know how to solve it

OpenStudy (er.mohd.amir):

since its select dark hair but its given that that has blue eye

OpenStudy (er.mohd.amir):

formula if select an event A but its given that event B then P(A/B)=P(A and B)/P(B)

OpenStudy (anonymous):

It asks for people who have dark hair and blue eyes. 55% have dark hair and 35% have dark hair and blue eyes. So it is 35 people out of 55. Thats why Er.Mohd.AMIR said 35 divided by 55. I hope you understand

OpenStudy (er.mohd.amir):

need more explanation

OpenStudy (anonymous):

OK thanks